# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 4.3 Additional measurement regions
Figure 4: The light cones of measured observables in canonical form are illustrated in a FDQC. The light cones of the logical operator and the measured observables $A_{j}$ attached thereto via Lemma 8 must intersect for teleportation to succeed.
We next turn to standard teleportation protocols that achieve a task distance $L>3vT$ (13), which exceeds the bound (43) of Lemma 11 for $M=2$. Theorem 3 Friedman et al. (2022a) suggests that this is possible if one uses additional measurements and accompanying feedback to link up the light cones of the additional measured observables attached to the logical operators, as depicted in Fig. 4.
As a starting point, consider a protocol $\mathcal{W}=\mathcal{W}_{m}\cdots\mathcal{W}_{1}$ in which $m$ teleportation protocols $\mathcal{W}_{s}$ are applied in series. Each protocol $\mathcal{W}_{s}$ satisfies $L_{s}=d(i_{s},f_{s})\leq v(3\,T_{s}-2)$ (43), where $f_{s}=i_{s+1}$ is the initial site of the subsequent protocol, so that the total distance (13) is
$L$ $=d(i_{1},f_{m})=\sum\limits_{s=1}^{m}L_{s}v\,\sum\limits_{s=1}^{m}(3\,T_{s}-2)$
@@
where $T=\sum_{s}T_{s}$ is the total depth. Unsurprisingly, applying protocols in series obeys the same bound (43).
However, an improvement can be realized by applying the protocols in parallel. Most of the channels in the protocol $\mathcal{W}_{s}$ do not overlap with those of $\mathcal{W}_{s-1}$. For convenience, suppose each protocol $\mathcal{W}_{s}$ has the same depth $T_{s}=T^{\prime}$. As explained following Lemma 11, each protocol $\mathcal{W}_{s}$ uses $M_{s}=2$ measurements to link a unitary “staircase” region of size $\ell_{s,{\rm uni}}\leq vT^{\prime}$ and another measurement-assisted “light-cone” region of size $\ell_{s,{\rm meas}}\leq 2v(T^{\prime}-1)$.
@@
$L_{\rm par}$ $\leq\frac{3}{4}v\,M\left(T-\frac{4}{3}\right)\,,~{}~{}$ (45)
@@
However, it is possible to exceed the bound (45) by constructing fully parallel protocols. In particular, by omitting the staircase segments of each protocol $\mathcal{W}_{s}$ except $\mathcal{W}_{1}$, the total depth can be cut in half compared to the parallelized case. As a result, the upper bound (45) is modified by sending $T\to 2T$ and subtracting $(m-1)vT$. The resulting bound for a generic standard teleportation protocol with $M>1$ is captured by Lemma 12.
@@
The proof of Lemma 12 appears in App. J. The bound (46) is saturated by parallel protocols in which $m=M/2$ pairs of measurements are spaced by a distance $\ell=2v(T-1)$ to daisy chain $m=\lfloor M/2\rfloor$ light-cone regions of size $\ell$ and a single staircase region of size $\ell_{0}\leq vT$. Note that Eq. 46 saturates the $k=1$ bound (30) Friedman et al. (2022a), up to $\mathsf{O}\left(1\right)$ offsets, while the teleportation bound (29) of Theorem 3 with $k=1$ is loose by a factor of two compared to the $k=1$ bound (46), as conjectured in Ref. 36.
At this point, we note that the “measurement regions” discussed in Ref. 36 are only unambiguously defined for optimal protocols. In that case, each pair of (adjacent) measurements used to daisy chain regions in the proof of Lemma 12 (see App. J) defines a distinct measurement region (the support of the pair of operators). There are $M=\lfloor M/2\rfloor$ such regions in total, where the $s$th region has size $v+1\leq\ell_{s}\leq 2v(T-1)$. However, we note that (i) the bounds herein are all phrased in terms of the number of measurement outcomes and (ii) an unambiguous delineation of measurement regions can be identified even in suboptimal protocols by considering string order, as we establish in Sec. 5.
@@ 4.4 Multiqubit teleportation
where all measured observables $A_{j}$ (32) commute with one another and with all final-state logical operators (7).
@@
The perspective is that a single region of size $\ell_{0}\leq vT$ can be daisy chained with $m$ regions of size $2k\leq\ell\leq 2v(T-1)$ using measurements of $2k$ sites. In the optimal case, the measurements are made on $2k$ consecutive sites, where the sites $12n-$ 1 and $2n$ correspond to the $n$th logical qubit. In this scenario, the optimal spacings of initial and final sites are given by $i_{n}=i_{1}+2(n-1)$ and $f_{n}=f_{1}+2(n-1)$, respectively.
Finally, we remark that the teleportation bound (47) is tighter than corresponding bound (29) of Theorem 3 and Ref. 36. Essentially, Eq. 29 is loose by a factor of two in the context of standard teleportation protocols, as initially conjectured in Ref. 36. While, in principle, it is possible that nonstandard teleportation protocols could exceed the teleportation bound (47)—and instead obey Eq. 29—we note that (i) both bounds have the same asymptotic scaling in terms of $L$, $M$, and $T$, which makes them equivalent in the context of such locality bounds, and (ii) as far as we are aware, the types of channels that we have eschewed in our analysis of optimal protocols in App. K—such as conditional measurements and non-Pauli recovery operations—cannot lead to violations of the standard bound (47). Hence, we expect that Eq. 47 optimally bounds all physical teleportation protocols.
@@ 5 The resource state is an SPT
We now derive the main result: Standard teleportation protocols involve preparing an SPT resource state $\lvert\Psi_{t}\rangle=\mathcal{U}\,\lvert\Psi_{0}\rangle$ (36) protected by a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry, measuring the corresponding string order parameters, and applying outcome-dependent Pauli gates. We briefly review SPTs in Sec. 5.1 before proving that the resource state $\lvert\Psi_{t}\rangle$ for $k=1$ teleportation has $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ string order. We then extend this result to $k>1$ in Sec. 5.3, where the protecting symmetry consists of (at least) $k$ copies of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, each corresponding to pairs of sublattices.
@@ 5.1 Overview of SPTs
meaning that applying the on-site symmetry operators $u_{j}(g)$ in $U(g)$ (49) to some subregion $\mathcal{I}$ only affects the symmetric state $\lvert\Psi\rangle$ near the boundaries of $\mathcal{I}$. Eq. (50) can be shown rigorously using matrix product states Pérez-García et al. (2008).
The operators $V_{L}(g)\,U_{\mathcal{I}}(g)\,V_{R}(g)$ in Eq. 50 are known as string order parameters Pollmann and Turner (2012). More generally, a string order parameter is any operator consisting of an on-site representation $u_{j}(g)$ (49) of $g\in\mathcal{G}$ in an interval $\mathcal{I}$ decorated by charged operators $V$ at the endpoints Pollmann and Turner (2012). Here, we can always choose endpoints such that the string order is perfect, meaning it has expectation value one. While the operators $u_{j}(g)$ in $U(g)$ (49) realize a linear representation of $g\in\mathcal{G}$, the endpoint operators $V_{L/R}(g)$ only realize a projective representation of $g\in\mathcal{G}$, i.e.,
@@
for all $g,h\in\mathcal{G}$, where $\omega(g,h)\in\mathsf{U}\left(1\right)$ is a complex phase. This phenomenon is known as symmetry fractionalization, and is the basis for classifying SPT phases in 1D Chen et al. (2011); Schuch et al. (2011); Pollmann et al. (2010).
@@
The fact that $\Omega(g,h)\neq 1$ for some $g,h\in\mathcal{G}$ is necessary and sufficient for the projective representation to belong to a nontrivial class is shown in Berkovich and Zhmud (1998). One can show that nontrivial string order (53) implies that (i) any gapped Hamiltonian with symmetry group $\mathcal{G}$ and ground state $\lvert\Psi\rangle$ has zero-energy edge modes Pollmann et al. (2010) and (ii) the state $\lvert\psi\rangle$ cannot be mapped to a product state via a FDQC consisting of local gates that respect the symmetry Huang and Chen (2015). Accordingly, nontrivial string order (53) of $\lvert\Psi\rangle$ is often regarded as synonymous with nontrivial SPT order of $\lvert\Psi\rangle$ (or the parent Hamiltonian). However, since SPT order delineates a phase of matter, it is only well defined in the thermodynamic limit Zeng and Wen (2015). Hence, establishing SPT order more carefully requires considering a family of states $\lvert\Psi\rangle$ defined on chains of increasing length, each of which possesses nontrivial string order in the sense of Def. 8.
5.2 Single-qubit case: $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$
We first prove that standard teleportation protocols $\mathcal{W}$ that teleport a single logical qubit ($k=1$), when written in canonical form (31), correspond to preparing an SPT resource state $\lvert\Psi_{t}\rangle$ (36), measuring its string order parameters, and applying Pauli recovery gates to the final site $f$. In particular, the nontrivial string order of $\lvert\Psi_{t}\rangle$ (36) is with respect to the discrete Abelian symmetry $\mathcal{G}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. The endpoint operators satisfy the Pauli algebra generated by $X$ and $Z$, which realize a projective representation of $\mathcal{G}$. As a first step toward proving that $\lvert\Psi_{t}\rangle$ (36) has nontrivial string order in the sense of Def. 8, we first present Lemma 15, establishing the existence of order parameters (53) when $\mathcal{I}$ is the full chain.
@@ 5.2 Single-qubit case: $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$
Suppose that the standard teleportation protocol $\mathcal{W}$ teleports $k=1$ logical qubit from $i$ to $f$ and has canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31). Then the resource state $\lvert\Psi_{t}\rangle$ (36) has string order characterized by, e.g., the order parameters
$\left\langle\Psi_{t}\middle|\,\mathcal{U}\,X_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{x}\,X_{f}\,\middle|\Psi_{t}\right\rangle$ $=1$ (54a)
$\left\langle\Psi_{t}\middle|\,\mathcal{U}\,Z_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{z}\,Z_{f}\,\middle|\Psi_{t}\right\rangle$ $=1\,,~{}~{}$ (54b)
@@
$\bar{A}_{\nu}\equiv\bar{A}_{1}^{\lambda_{1,\nu}}\cdots\bar{A}_{M}^{\lambda_{M,\nu}}\,,~{}~{}$ (55)
i.e., the product of the involutory parts (4) of all necessary measured observables $\bar{A}_{j}$ (see Cor. 14) attached to $\sigma^{\nu}_{f}$ by Lemma 8. Hence, the string order parameters (15) commute in the bulk region $(i+vT,f)$, i.e.,
$\left[\bar{A}_{x},\,\bar{A}_{z}\right]=0\,,~{}~{}$ (56)
@@
The proof of Lemma 15 appears in App. M, and follows straightforwardly from previous results. The end-to-end string order parameters are depicted graphically in Fig. 5(a). The operators above $\mathcal{U}$ represent the two string order parameters (15) acting on the resource state $\lvert\Psi_{t}\rangle$ (36), with the anticommuting endpoint operators color coded. In the Heisenberg picture, the order parameters evolve under $\mathcal{U}$ into the stabilizers $\mathsf{S}_{x,z}\in\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$ at the bottom of Fig. 5(a), which have expectation value one (50) acting on the initial state $\lvert\Psi_{0}\rangle$ (6).
Figure 5: The construction of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ string order parameters (a) on the endpoints and (b) in the bulk, where the colored operators represent the anticommuting endpoints.
However, in general, a state is only said to have string order (53) if a similar relation holds for many choices of endpoints. As it turns out, the string order parameters (15) in Lemma 15, with right endpoint $f$ and left endpoint in $[i,i+vT]$, are not the only valid choice for the resource state $\lvert\Psi_{t}\rangle$ (36). In Lemma 16, we prove that a set of valid string order parameters (53) corresponding to various subintervals $\mathcal{I}\subset[i,f]$ can be identified.
Lemma 16 ($\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ string order).
In addition to the end-to-end string order parameter (15) of Lemma 15, the resource state $\lvert\Psi_{t}\rangle$ (36) exhibits nontrivial string order (see Def. 8) captured by the order parameters
$\mathcal{S}_{\nu,a,b}=\left(\mathcal{U}\,\Sigma^{\nu}_{a}\,\mathcal{U}^{\dagger}\right)\,\bar{A}_{\nu,a+1}\cdots\bar{A}_{\nu,b}\,\left(\mathcal{U}\,\Sigma^{\nu}_{b}\,\mathcal{U}^{\dagger}\right)\,,~{}~{}$ (57)
@@
$\Sigma^{\nu}_{s}$ $=\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu,1}\cdots\bar{A}_{\nu,s}\,\mathcal{U}=\Sigma^{\nu}_{s-1}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu,s}\,\mathcal{U}\,,~{}~{}$ (59)
where $\Sigma^{\nu}_{0}=\sigma^{\nu}_{i}$ and $\left\{\Sigma^{x}_{s},\,\Sigma^{z}_{s}\right\}=0$, as required. The $m$ operators $\bar{A}_{s}$ correspond to products of the measured observables $\bar{A}_{j}$ (4) in disjoint regions labeled $s$, such that $\mathcal{U}^{\dagger}\,\bar{A}_{s+1}\,\mathcal{U}$ has no overlap with $\Sigma^{\nu}_{s}$ and $\Sigma^{\nu}_{s}$ acts to the right of $\Sigma^{\nu}_{s-1}$ and to the left of $\Sigma^{\nu}_{s+1}$.
Since, for all $a<b\in[1,m]$, $\lvert\Psi_{t}\rangle$ (36) satisfies
$\left\langle\Psi_{t}\middle|\mathcal{S}_{\nu,a,b}\middle|\Psi_{t}\right\rangle=1\,,~{}~{}$ (60)

### v3

@@ 4.3 Additional measurement regions
Figure 4: The light cones of measured observables in canonical form are illustrated in a FDQC. The light cones of the logical operator and the measured observables $A^{\,}_{j}$ attached thereto via Lemma 8 must intersect for teleportation to succeed.
We next turn to standard teleportation protocols that achieve a task distance $L>3vT$ (13), which exceeds the bound (43) of Lemma 11 for $M=2$. Theorem 3 [42] suggests that this is possible if one uses additional measurements and accompanying feedback to link up the light cones of the additional measured observables attached to the logical operators, as depicted in Fig. 4.
As a starting point, consider a protocol $\mathcal{W}=\mathcal{W}^{\,}_{m}\cdots\mathcal{W}^{\,}_{1}$ in which $m$ teleportation protocols $\mathcal{W}^{\,}_{s}$ are applied in series. Each protocol $\mathcal{W}^{\,}_{s}$ satisfies $L^{\,}_{s}=d(i_{s},f_{s})\leq v(3\,T^{\,}_{s}-2)$ (43), where $f_{s}=i_{s+1}$ is the initial site of the subsequent protocol, so that the total distance (13) is
$L$ $=d(i^{\,}_{1},f^{\,}_{m})=\sum\limits_{s=1}^{m}L^{\vphantom{\prime}}_{s}v\,\sum\limits_{s=1}^{m}(3\,T^{\vphantom{\prime}}_{s}-2)$
@@
where $T=\sum_{s}T^{\,}_{s}$ is the total depth. Unsurprisingly, applying protocols in series obeys the same bound (43).
However, an improvement can be realized by applying the protocols in parallel. Most of the channels in the protocol $\mathcal{W}^{\,}_{s}$ do not overlap with those of $\mathcal{W}^{\,}_{s-1}$. For convenience, suppose each protocol $\mathcal{W}^{\,}_{s}$ has the same depth $T^{\,}_{s}=T^{\prime}$. As explained following Lemma 11, each protocol $\mathcal{W}^{\,}_{s}$ uses $M^{\,}_{s}=2$ measurements to link a unitary “staircase” region of size $\ell^{\,}_{s,{\rm uni}}\leq vT^{\prime}$ and another measurement-assisted “light-cone” region of size $\ell^{\,}_{s,{\rm meas}}\leq 2v(T^{\prime}-1)$.
@@
$L^{\vphantom{\prime}}_{\rm par}$ $\leq\frac{3}{4}v\,M\left(T-\frac{4}{3}\right)\,,~{}~{}$ (45)
@@
However, it is possible to exceed the bound (45) by constructing fully parallel protocols. In particular, by omitting the staircase segments of each protocol $\mathcal{W}^{\,}_{s}$ except $\mathcal{W}^{\,}_{1}$, the total depth can be cut in half compared to the parallelized case. As a result, the upper bound (45) is modified by sending $T\to 2T$ and subtracting $(m-1)vT$. The resulting bound for a generic standard teleportation protocol with $M>1$ is captured by Lemma 12.
@@
The proof of Lemma 12 appears in App. J. The bound (46) is saturated by parallel protocols in which $m=M/2$ pairs of measurements are spaced by a distance $\ell=2v(T-1)$ to daisy chain $m=\lfloor M/2\rfloor$ light-cone regions of size $\ell$ and a single staircase region of size $\ell^{\,}_{0}\leq vT$. Note that Eq. 46 saturates the $k=1$ bound (30) [42], up to $\mathsf{O}\left(1\right)$ offsets, while the teleportation bound (29) of Theorem 3 with $k=1$ is loose by a factor of two compared to the $k=1$ bound (46), as conjectured in Ref. 42.
At this point, we note that the “measurement regions” discussed in Ref. 42 are only unambiguously defined for optimal protocols. In that case, each pair of (adjacent) measurements used to daisy chain regions in the proof of Lemma 12 (see App. J) defines a distinct measurement region (the support of the pair of operators). There are $M=\lfloor M/2\rfloor$ such regions in total, where the $s$th region has size $v+1\leq\ell^{\,}_{s}\leq 2v(T-1)$. However, we note that (i) the bounds herein are all phrased in terms of the number of measurement outcomes and (ii) an unambiguous delineation of measurement regions can be identified even in suboptimal protocols by considering string order, as we establish in Sec. 5.
@@ 4.4 Multiqubit teleportation
where all measured observables $A^{\,}_{j}$ (32) commute with one another and with all final-state logical operators (7).
@@
The perspective is that a single region of size $\ell^{\,}_{0}\leq vT$ can be daisy chained with $m$ regions of size $2k\leq\ell\leq 2v(T-1)$ using measurements of $2k$ sites. In the optimal case, the measurements are made on $2k$ consecutive sites, where the sites $12n-$ 1 and $2n$ correspond to the $n$th logical qubit. In this scenario, the optimal spacings of initial and final sites are given by $i_{n}=i_{1}+2(n-1)$ and $f_{n}=f_{1}+2(n-1)$, respectively.
Finally, we remark that the teleportation bound (47) is tighter than corresponding bound (29) of Theorem 3 and Ref. 42. Essentially, Eq. 29 is loose by a factor of two in the context of standard teleportation protocols, as initially conjectured in Ref. 42. While, in principle, it is possible that nonstandard teleportation protocols could exceed the teleportation bound (47)—and instead obey Eq. 29—we note that (i) both bounds have the same asymptotic scaling in terms of $L$, $M$, and $T$, which makes them equivalent in the context of such locality bounds, and (ii) as far as we are aware, the types of channels that we have eschewed in our analysis of optimal protocols in App. K—such as conditional measurements and non-Pauli recovery operations—cannot lead to violations of the standard bound (47). Hence, we expect that Eq. 47 optimally bounds all physical teleportation protocols.
@@ 5 The resource state is an SPT
We now derive the main result: Standard teleportation protocols involve preparing an SPT resource state $\lvert\Psi_{t}\rangle=\mathcal{U}\,\lvert\Psi_{0}\rangle$ (36) protected by a $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ symmetry, measuring the corresponding string order parameters, and applying outcome-dependent Pauli gates. We briefly review SPTs in Sec. 5.1 before proving that the resource state $\lvert\Psi_{t}\rangle$ for $k=1$ teleportation has $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ string order. We then extend this result to $k>1$ in Sec. 5.3, where the protecting symmetry consists of (at least) $k$ copies of $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$, each corresponding to pairs of sublattices.
We stress that SPT phases are distinct from phases with “instrinsic” topological order like the toric code [67]. In some sense, the SPT phase can be viewed as a “trivial” topological order with no anyon content (i.e., a only a single, trivial anyonic excitation). Another important distinction is that SPT states are not long-range entangled; as a result, they can be prepared using FDQCs [31], making them promising resource states $\lvert\Psi_{t}\rangle$ (36). Additionally, while it is possible to teleport using a topologically ordered resource state—e.g., the ground state of the toric code [68]—it is inefficient, because preparing the long-range-entangled resource state requires the same resources as teleportation itself [42].
@@ 5.1 Overview of SPTs
meaning that applying the on-site symmetry operators $u_{j}(g)$ in $U(g)$ (49) to some subregion $\mathcal{I}$ only affects the symmetric state $\lvert\Psi\rangle$ near the boundaries of $\mathcal{I}$. Eq. 50 can be shown rigorously using matrix product states [44].
The operators $V_{L}(g)\,U_{\mathcal{I}}(g)\,V_{R}(g)$ in Eq. 50 are known as string order parameters [38]. More generally, a string order parameter is any operator consisting of an on-site representation $u_{j}(g)$ (49) of $g\in\mathcal{G}$ in an interval $\mathcal{I}$ decorated by charged operators $V$ at the endpoints [38]. Here, we can always choose endpoints such that the string order is perfect, meaning it has expectation value one. While the operators $u_{j}(g)$ in $U(g)$ (49) realize a linear representation of $g\in\mathcal{G}$, the endpoint operators $V_{L/R}(g)$ only realize a projective representation of $g\in\mathcal{G}$, i.e.,
@@
for all $g,h\in\mathcal{G}$, where $\omega(g,h)\in\mathsf{U}\left(1\right)$ is a complex phase. This phenomenon is known as symmetry fractionalization, and is the basis for classifying SPT phases in 1D [31, 33, 32].
@@
The fact that $\Omega(g,h)\neq 1$ for some $g,h\in\mathcal{G}$ is necessary and sufficient for the projective representation to belong to a nontrivial class is shown in [69]. One can show that nontrivial string order (53) implies that (i) any gapped Hamiltonian with symmetry group $\mathcal{G}$ and ground state $\lvert\Psi\rangle$ has zero-energy edge modes [32] and (ii) the state $\lvert\psi\rangle$ cannot be mapped to a product state via a FDQC consisting of local gates that respect the symmetry [70]. Accordingly, nontrivial string order (53) of $\lvert\Psi\rangle$ is often regarded as synonymous with nontrivial SPT order of $\lvert\Psi\rangle$ (or the parent Hamiltonian). However, since SPT order delineates a phase of matter, it is only well defined in the thermodynamic limit [71]. Hence, establishing SPT order more carefully requires considering a family of states $\lvert\Psi\rangle$ defined on chains of increasing length, each of which possesses nontrivial string order in the sense of Def. 8.
5.2 Single-qubit case: $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$
We first prove that standard teleportation protocols $\mathcal{W}$ that teleport a single logical qubit ($k=1$), when written in canonical form (31), correspond to preparing an SPT resource state $\lvert\Psi^{\,}_{t}\rangle$ (36), measuring its string order parameters, and applying Pauli recovery gates to the final site $f$. In particular, the nontrivial string order of $\lvert\Psi^{\,}_{t}\rangle$ (36) is with respect to the discrete Abelian symmetry $\mathcal{G}=\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$. The endpoint operators satisfy the Pauli algebra generated by $X^{\vphantom{*}}$ and $Z^{\vphantom{*}}$, which realize a projective representation of $\mathcal{G}$. As a first step toward proving that $\lvert\Psi^{\,}_{t}\rangle$ (36) has nontrivial string order in the sense of Def. 8, we first present Lemma 15, establishing the existence of order parameters (53) when $\mathcal{I}$ is the full chain.
@@ 5.2 Single-qubit case: $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$
Suppose that the standard teleportation protocol $\mathcal{W}$ teleports $k=1$ logical qubit from $i$ to $f$ and has canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31). Then the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) has string order characterized by, e.g., the order parameters
$\left\langle\Psi^{\vphantom{\prime}}_{t}\middle|\,\mathcal{U}\,X^{\vphantom{*}}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{x}\,X^{\vphantom{*}}_{f}\,\middle|\Psi^{\vphantom{\prime}}_{t}\right\rangle$ $=1$ (54a)
$\left\langle\Psi^{\vphantom{\prime}}_{t}\middle|\,\mathcal{U}\,Z^{\vphantom{*}}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{z}\,Z^{\vphantom{*}}_{f}\,\middle|\Psi^{\vphantom{\prime}}_{t}\right\rangle$ $=1\,,~{}~{}$ (54b)
@@
$\bar{A}^{\vphantom{\prime}}_{\nu}\equiv\bar{A}_{1}^{\lambda^{\,}_{1,\nu}}\cdots\bar{A}_{M}^{\lambda^{\,}_{M,\nu}}\,,~{}~{}$ (55)
i.e., the product of the involutory parts (33) of all necessary measured observables $\bar{A}^{\,}_{j}$ (see Cor. 14) attached to $\sigma^{\nu}_{f}$ by Lemma 8. Hence, the string order parameters (54) commute in the bulk region $(i+vT,f)$, i.e.,
$\left[\bar{A}^{\vphantom{\prime}}_{x},\,\bar{A}^{\vphantom{\prime}}_{z}\right]=0\,,~{}~{}$ (56)
@@
The proof of Lemma 15 appears in App. M, and follows straightforwardly from previous results. The end-to-end string order parameters are depicted graphically in Fig. 5(a). The operators above $\mathcal{U}$ represent the two string order parameters (54) acting on the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36), with the anticommuting endpoint operators color coded. In the Heisenberg picture, the order parameters evolve under $\mathcal{U}$ into the stabilizers $\mathsf{S}^{\,}_{x,z}\in\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$ at the bottom of Fig. 5(a), which have expectation value one (50) acting on the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6).
Figure 5: The construction of $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ string order parameters (a) on the endpoints and (b) in the bulk, where the colored operators represent the anticommuting endpoints.
However, in general, a state is only said to have string order (53) if a similar relation holds for many choices of endpoints. As it turns out, the string order parameters (54) in Lemma 15, with right endpoint $f$ and left endpoint in $[i,i+vT]$, are not the only valid choice for the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36). In Lemma 16, we prove that a set of valid string order parameters (53) corresponding to various subintervals $\mathcal{I}\subset[i,f]$ can be identified.
Lemma 16 ($\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ string order).
In addition to the end-to-end string order parameter (54) of Lemma 15, the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) exhibits nontrivial string order (see Def. 8) captured by the order parameters
$\mathcal{S}^{\vphantom{\prime}}_{\nu,a,b}=\left(\mathcal{U}\,\Sigma^{\nu}_{a}\,\mathcal{U}^{\dagger}\right)\,\bar{A}^{\vphantom{\prime}}_{\nu,a+1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,b}\,\left(\mathcal{U}\,\Sigma^{\nu}_{b}\,\mathcal{U}^{\dagger}\right)\,,~{}~{}$ (57)
@@
$\Sigma^{\nu}_{s}$ $=\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,s}\,\mathcal{U}=\Sigma^{\nu}_{s-1}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,s}\,\mathcal{U}\,,~{}~{}$ (59)
where $\Sigma^{\nu}_{0}=\sigma^{\nu}_{i}$ and $\left\{\Sigma^{x}_{s},\,\Sigma^{z}_{s}\right\}=0$, as required. The $m$ operators $\bar{A}^{\,}_{s}$ correspond to products of the measured observables $\bar{A}^{\,}_{j}$ (33) in disjoint regions labeled $s$, such that $\mathcal{U}^{\dagger}\,\bar{A}^{\,}_{s+1}\,\mathcal{U}$ has no overlap with $\Sigma^{\nu}_{s}$ and $\Sigma^{\nu}_{s}$ acts to the right of $\Sigma^{\nu}_{s-1}$ and to the left of $\Sigma^{\nu}_{s+1}$.
Since, for all $a<b\in[1,m]$, $\lvert\Psi^{\,}_{t}\rangle$ (36) satisfies
$\left\langle\Psi^{\vphantom{\prime}}_{t}\middle|\mathcal{S}^{\vphantom{\prime}}_{\nu,a,b}\middle|\Psi^{\vphantom{\prime}}_{t}\right\rangle=1\,,~{}~{}$ (60)